ELECTRON SPIN RESONANCE SPECTROSCOPY STUDIES OF RADICAL IONS TRAPPED IN PARA-HYDROGEN MATRICES

\underline{WILLIAM DEREK BUSH}, MATTHEW S. WARD and LON B. KNIGHT, JR., Department of Chemistry, Furman University, Greenville, SC, 29613.

The ability to trap atomic cation radicals in para-hydrogen at 2K for ESR (electron spin resonance) studies has been demonstrated for the first time in a series of recent experiments in our laboratory. Comparative magnetic parameters (g and A tensors) for ^{2}^{5}Mg^{+}, ^{6}^{7}Zn^{+} and ^{1}^{9}^{9}^{,}^{2}^{0}^{1}Hg^{+} have been measured in neon, argon, krypton, and p-H_{2} matrices. The cations were generated by pulsed laser ablation of the metals, x-irradiation or by a neon discharge photoionization source at 16.8 eV. Atomic hydrogen lines were observed to have unusually narrow linewidths in p-H_{2}, in agreement with earlier findings reported by other investigators. The relative intensity of satellite transitions seems to reflect the degree of conversion from n-H_{2} to p-H_{2}.